OpenStudy (anonymous):

How can you identify the effect on the graph of replacing f(x) by f(x) + k, k f(x), f(kx), and f(x + k) for specific values of k? @gabylovesu

OpenStudy (anonymous):

f(x)+k is when you go up on the graph f(x+k) means that you graph will shift l left

OpenStudy (anonymous):

okay and can you help with another?

OpenStudy (anonymous):

How can you compare properties of two functions each represented in a different way? @gabylovesu

OpenStudy (anonymous):

can you help @Kainui

OpenStudy (kainui):

You should get a graphing calculator or use a website like www.wolframalpha.com to inspect these yourself. So compare the graph of \[\Large f(x)=x^2 \] with \[\Large f(x+3)=(x+3)^2\] to see for yourself what's going on.